OpenStudy (anonymous):

Car A began a journey from a point at 9 am, traveling at 30 mph. At 10 am car B started traveling from the same point at 40 mph in the same direction as car A. At what time will car B pass car A?

OpenStudy (anonymous):

Let after t hours the distances D1 traveled by car A => D1 = 30 t Car B starts at 10 am and will therefore have spent one hour less than car A when it passes it. After (t - 1) hours, distance D2 traveled by car B => D2 = 40 (t - 1) When car B passes car A, they are at the same distance from the starting point and therefore D1 = D2 => 30 t = 40 (t - 1) Solve the equation for t, => 30 t = 40t - 40 => 10 t = 40 => t = 4 => Car B passes car A at = 9 + 4 = 13 pm.

OpenStudy (anonymous):

that should help out :D